OpenStudy (anonymous):

you spin a spinner and the probability that the spinner will land on red is 18%. Suppose you spin the spinner 4 times. What is the probability that the spinner will land on red at least 2 times?

OpenStudy (kropot72):

This can be solved by the binomial distribution: \[P(0\ red)=0.82^{4}=0.4521\] \[P(1\ red)=4\times 0.18\times 0.82^{3}=0.3970\] P(0 red) + P(1 red) = 0.8491 The probability that the spinner will land on red at least 2 times is 1.0000 - 0.8491 = ?

OpenStudy (anonymous):

your answer is 15.1% but that's not a choice: 10.1%, 11.9%; 24.3%; or 63.8%

OpenStudy (anonymous):

do u see the choices

OpenStudy (kropot72):

Can you check that the question has been posted correctly. I am confident in the solution with the question as it is now.